Circular No. 6638

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) A. Fitzsimmons, Queen’s University, Belfast; and G. Cremonese, As-

tronomical Observatory, Padua, on behalf of the European Hale-Bopp Team, report on further observations of the tails of comet C/1995 O1 (cf. IAUC 6631, 6634): “The 4.2-m William Herschel Telescope (+ Utrecht Echelle Spectrograph) at Observatorio del Roque de los Muchachos was used on Apr. 23.9 and 24.9 UT to obtain spectra with a 4′.5 × 1′′.1 slit at various positions on the dust tail up to 2^{o} from the nucleus. Broad sodium emission lines were seen at all locations, with a blue edge corresponding to the geocentric velocity of the comet and a width of typically ∼ 50 km/s. These observations imply that dust grains are a major source of sodium in this comet. We also report the detection with the same instrument of neutral potassium (769.9 nm) emission associated with the comet in spectra obtained on the nucleus and at a distance of 10′′ on Apr. 20.9.”

V1333 AQUILAE S. A. Ilovaisky and C. Chevalier, Observatoire de Haute-Provence,

communicate: “Subsequent to the x-ray detection of Aquila X-1 with the XTE All Sky Monitor around mid-Jan. (IAUC 6558), the optical counterpart of this recurrent transient was observed with the 1.2-m telescope on Feb. 20 and found to be bright. Images taken on Mar. 30 show that the object had already faded to quiescence. The XTE ASM quick-look x-ray lightcurve, publicly available on the Web at http://space.mit.edu/XTE/ asmlc/srcs/ASM.html, shows that the source had turned off by early Mar. During the previous activity cycle, the optical object was observed at Haute-Provence between 1996 Feb. 28 and Oct. 12. Optical turn-on was detected on 1996 May 29 (IAUC 6416) and optical turn-off on Aug. 19. Both of these dates agree closely (within one or two days) with those derived from the ASM data for that low-intensity event.”

SUPERNOVAE 1997bp, 1997bq, AND 1997bs Unfiltered CCD magnitude estimates by S. Moretti and S. Tomaselli,

Forlí, Italy, with a 0.35-m telescope (reference stars from GSC): SN 1997 bp in NGC 4680, Apr. 23.86 UT, 15.0; SN 1997bq in NGC 3147, Apr. 18.92, 14.8; SN 1997bs in NGC 3627, Apr. 23.89, 17.4.

1997 April 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green